OpenStudy (anonymous):

What is the next number in the sequence? 60, 28, 12, 4,

OpenStudy (isaiah):

i think the answer might be zero i could be wrong so dont listen to me lol

OpenStudy (amistre64):

its at least less than 4 :)

OpenStudy (amistre64):

what techniques do you know of to help out?

OpenStudy (anonymous):

the ansrews are 2, 0, -4, -8

OpenStudy (amistre64):

1/2 -2

OpenStudy (amistre64):

60/2 - 2 = 28 28/2 - 2 = 12 12/2 - 2 = 4 4/2 - 2 = 0

OpenStudy (anonymous):

sweet, thanks

OpenStudy (amistre64):

other than trial and error; i got no good way of doing these :)

OpenStudy (anonymous):

What is the scientific notation for the number of kilometers light will travel in 60 seconds? (time 1sec) (distance 186,000mi /300,000 km)

OpenStudy (anonymous):

I used straight subtraction but ended up with the same answer: \[60-28=32\] \[28-12=16\] \[12-4=8\] Reduction of a power of 2 for the difference each step so: \[4-4=0\] Similar answer.